OpenStudy (anonymous):

If 20 lb of rice and 10 lb of potatoes cost $16.20 and 30 lb of rice and 12 lb of potatoes cost $23.04, how much will 10 lb of rice and 50 lbs of potatoes cost?

OpenStudy (mathstudent55):

Let x = price of rice per pound Let y = price of potaotes per pound 20x + 10y = 16.2 30x + 12y = 23.04 Solve the system of equations for x and y. Then substitute x and y in 10x + 50y and evaluate the expression.

OpenStudy (anonymous):

Elimination is faster

OpenStudy (anonymous):

The way @mathstudent55 has it is the way i need it set up.. but idk how to do elimination anyway 20x+10y=16.2 R1 30x+50y=23.04 R2

OpenStudy (mathstudent55):

To use elimination you need to add the two equations and eliminate one variable. Sometimes you may have a system of equations in which for, example, one equation has 2x and the other one has -2x. Simply by adding the eqquatios, you eliminate x and you just need to solve for y. Most of the time, though, you need to multiply one equation or both by some number to be able to eliminate a variable. Let's take this example. The first equation has 20x and the second equation has 30x. If you had -60y in the first equation and 60y in the second equation, the y's would add up to zero. In order to do that, let's multiply the entire first equation by -3 and let's multiply the entire second equation by 2. Then we'll simplify both equations and write one above the other. (-3)(20x) + (-3)(10y) = (-3)(16.2) -60x - 30y = -48.6 (2)(30x) + (2)(12y) = (2)(23.04) 60x + 24y = 46.08 -60x - 30y = -48.6 60x + 24y = 46.08 --------------------(add the two equations) -6y = -2.52 Divide both sides by -6: -6y/(-6) = -2.52/(-6) y = 0.42 Now substitute y = 0.42 in either of the two original equations: Let's use th efirst one 20x + 10y = 16.2 20x + 10(0.42) = 16.2 20x + 4.2 = 16.2 20x = 12 x = 0.60 Now that you know the prices per pound of rice and potatoes, use the prices in 10x + 50y to evaluate what the question asks for.